# Robustness-Aware 3D Object Detection in Autonomous Driving: A Review and Outlook

### v1

@@ III-D1 Accuracy
Accuracy is a focal point of interest in the majority of research articles and reviews and is indeed of paramount importance. While accuracy can be reflected through AP (average precision), considering AP alone for comparison may not provide a comprehensive view, as different methodologies may exhibit substantial differences due to differing paradigms. As shown in Fig. 5 (a), we selected 10 representative methods (including classic and latest research) for comparison, and it is evident that there are significant metric disparities between monocular 3D object detection [21, 54, 305, 18, 55, 56, 76, 11, 75, 301] and stereo-based 3D object detection [358, 310, 359, 24, 360, 13, 361, 362, 28, 363]. The current scenario indicates that the accuracy of monocular 3D object detection is far lower than that of stereo-based 3D object detection. Stereo-based 3D object detection leverages the capture of images from two different perspectives of the same scene to obtain depth information. The greater the baseline between cameras, the wider the range of depth information captured. As shown in Fig. 5 (b), there existed monocular 3D object detection methods [364, 7, 67, 69, 65] on the nuScenes dataset 2021 ago, but no related research on Stereo-based 3D object detection. Starting from 2021, monocular methods have gradually been supplanted by multi-view (bird’s-eye-view perception) 3D object detection Multi-view [88, 89, 90, 26, 365, 330, 25, 329, 336, 337], leading to a significant improvement in mAP. The emergence of the novel bird’s-eye-view paradigm and the increase in sensor quantity have had a substantial impact on mAP. It can be observed that initially the disparity between DD3D [69] and DETR3D [88] was not prominent, but with the continuous enhancement of multi-view 3D object detection, particularly with the advent of novel works such as Far3D [337], the gap has widened. In other words, at present, Camera-xonly 3D object detection methods on multi-camera datasets like nuScenes [256] are predominantly based on bird’s-eye-view perception. If we consider accuracy solely from this single dimension, the increase in sensor quantity has led to a significant improvement in accuracy metrics (including mAP, NDS, AP, etc.).
@@ III-D2 Latency
Latency holds paramount importance in the realm of autonomous driving. It refers to the time required for a system to react to input signals, encompassing the entire process from sensor data acquisition to system decision-making and execution of actions. In autonomous driving, stringent requirements are imposed on latency, as any form of delay can lead to severe consequences. The following aspects underscore the importance of latency in autonomous driving.
Real-time responsiveness Autonomous driving systems need to demonstrate exceptional real-time responsiveness. Timely decision-making and actions are crucial for collision avoidance, adapting to traffic changes, and ensuring vehicle safety.
Safety High latency may result in the system’s inability to timely detect and respond to potential hazardous situations. Timely responses are a key factor in ensuring driving safety.
User Experience For passengers and other road users, a smooth and coherent driving experience is crucial. High latency may lead to an uncomfortable driving experience or even induce anxiety.
Interactivity Autonomous vehicles need to interact with other vehicles, pedestrians, and infrastructure. Low latency ensures timely communication and coordination, thereby enhancing the overall efficiency of the transportation system.
Emergency Response In emergency situations, such as the sudden appearance of obstacles or rapid changes in traffic conditions, the system needs to react quickly to mitigate potential dangers.
In the field of 3D object detection, Latency (Frames Per Second, FPS) and Accuracy are critical metrics for assessing the performance of algorithms. As shown in Fig. 6 (a), the chart for monocular and stereo 3D object detection illustrates the relationship between Average Precision (AP) at the moderate difficulty level of the KITTI dataset and FPS. Fig. 6 (b) shows the relationship between the nuScenes Detection Score (NDS) and FPS for monocular and multi-view 3D object detection. These FPS were obtained using a NVIDIA A100 graphics card, while the performance metrics AP and NDS are derived from the original papers.
Specifically, monocular-based 3D object detection, relying on data from a single camera, typically has lower computational requirements, thus achieving a higher FPS. However, due to the absence of depth information, its accuracy is often inferior to that of stereo or multi-view systems. Stereo-based 3D object detection, utilizing disparity information from images captured by dual cameras, enhances the accuracy of depth estimation but also introduces greater computational complexity, which may reduce FPS. Multi-view detection merges data from several cameras to provide richer scene information, which further improves accuracy. This method requires more extensive data processing, hence demanding greater computational power and algorithmic optimization to sustain a reasonable FPS level. Notably, there are no stereo-based 3D object detection methods represented on the nuScenes, with the monocular method FCOS3D[65] being particularly emblematic as it was introduced in 2021. With time and optimization, multi-view 3D object detection has rapidly developed in terms of both accuracy and latency.
In conclusion, for the realization of safe autonomous driving, 3D object detection algorithms must balance between Latency and Accuracy. While monocular detection is fast, it lacks precision, conversely, stereo and multi-view methods are accurate but slower. Future research should not only maintain high precision but also place greater emphasis on increasing FPS and reducing Latency to meet the dual requirements of real-time responsiveness and safety in autonomous driving.
@@ III-D3 Robustness
Robustness constitutes a pivotal factor in the safety perception of autonomous driving, representing a topic of significant attention that has been previously overlooked in comprehensive reviews. In the current meticulously designed clean datasets and benchmarks such as KITTI [255], nuScenes [256], and Waymo [262], this aspect is not commonly addressed. Presently, research works[253, 254, 388, 287, 288, 389, 390] like RoboBEV[253], Robo3D[288] on 3D object detection incorporate considerations of robustness, exemplified by factors such as sensor misses, as illustrated in Fig. 7. They have adopted a methodology involving the introduction of disturbances in datasets relevant to 3D object detection to assess robustness. This includes the introduction of various types of noise, such as variations in weather conditions, sensor malfunctions, motion disturbances, and object-related perturbations, aimed at unraveling the distinct impacts of different noise sources on the model. Typically, most papers investigating robustness conduct evaluations by introducing noise to the validation sets of clean datasets, such as KITTI [255], nuScenes [256], and Waymo [262]. Additionally, we highlight findings from Ref. [254], where KITTI-C[254], and nuScenes-C[254], are emphasized as examples to illustrate the results of Camera-Only 3D object detection methods. Tables IV and V provide an overall comparison, revealing that, in general, Camera-Only methods are less robust compared to LiDAR-Only and multi-model fusion methods. They are highly susceptible to various types of noise. In KITTI-C, three representative works—SMOKE[21], PGD[67], and ImVoxelNet[391]—show consistently lower overall performance and reduced robustness to noise. In nuScenes-C, noteworthy methods such as DETR3D[88] and BEVFormer[26] exhibit greater robustness compared to FCOS3D [65] and PGD[67], suggesting that as the number of sensors increases, overall robustness improves. In conclusion, future Camera-Only methods need to consider not only cost factors and Accuracy metrics (mAP, NDS, etc.) but also factors related to safety perception and robustness. Our analysis aims to provide valuable insights for the safety of future autonomous driving systems.
| Corruptions | LiDAR-Only | Camera-Only | multimodal | |||||||||||
| SECOND | PointPillars | PointRCNN | PV-RCNN | Part-A | 3DSSD | SMOKE | PGD | ImVoxelNet | EPNet | Focals Conv | LoGoNet * | VirConv-S * | ||
@@
| Corruptions | LiDAR-Only | Camera-Only | multimodal | |||||||||
| PointPillars | SSN | CenterPoint | FCOS3D | PGD | DETR3D | BEVFormer | FUTR3D | TransFusion | BEVFusion | D.I.* | ||
@@
IV LiDAR-based 3D Object Detection
In this section, we introduce LiDAR-based 3D object detection which utilizes point clouds as input data and extracts point cloud features to predict 3D objects. The point cloud (PC) is a set of points in Euclidean space, which can be expressed as , where 3 and C are the coordinate (x, y, z-axis) and extra feature (reflection intensity in general) respectively, and N is the number of points. The characteristics of the points are the following: (1) Sparsity. The point clouds are formally distributed discretely, and the location and distance of each point are irregularly distributed in 3D space. (2) Invariance. For example, when different sensors find an object, the order in which the points are collected is disparate. It poses that the network is hard to fit well because the points are all different. Therefore, putting CNN networks on point clouds causes information loss and makes it hard to train, and it usually requires preprocessing operations on point clouds to solve such problems.
Compared to camera-based methods, LiDAR captures precise 3D information, enabling LiDAR-based approaches to achieve higher detection accuracy and robustness, especially in extreme weather conditions [254]. Because in comparison to optical radiation, the laser beams emitted by LiDAR systems can penetrate certain weather disturbances, such as raindrops and haze, with slight interference. However, the high cost of LiDAR remains one of the main barriers to large-scale adoption of LiDAR-based methods. In addition, the lack of semantic information leads to poor classification performance of LiDAR-based methods. Generally, as shown in Fig. 8, LiDAR-based methods can be categorized into four types: (1) view-based 3D object detection, (2) voxel-based 3D object detection, (3) point-based 3D object detection, (4) point-voxel-based 3D object detection. In contrast to previous reviews [1, 344, 345, 346, 347, 242, 2, 348, 349, 350, 269, 252, 270, 271, 272, 273, 274, 275], our survey extends beyond the conventional classifications of LiDAR-based methods. We adopt a more foundational idea to class LiDAR-based methods based on their core data representations (such as 2D Bird’s Eye View (BEV), Voxel, Pillar, etc.) and underlying model structure (base models including Convolutional Neural Networks (CNN), Transformers, PointNet, and others). This methodological restructuring aims to provide a comprehensive understanding of the technological paradigms at the heart of LiDAR-based methods, analyzing and classifying these systems from a more essential, technical lineage perspective.
@@ IV-A View-based 3D object detection
The core idea of the view-based methods is to transform point clouds into pseudo-image representations from a bird’s-eye view (BEV) or range view by plane [392], cylindrical [393], or spherical [394] projections. In these representations, each pixel contains 3D spatial information rather than RGB values. Due to the dense representation of pseudo-images, traditional or specialized 2D convolutions can be seamlessly applied to range images, making the feature extraction process highly efficient. However, compared to other LiDAR-based methods, detection using range views is more susceptible to occlusion and scale variations. Hence, some methods [175, 173] transform the data representation from the range view to the bird’s eye view (BEV). Furthermore, due to the projection of distant 3D spatial points becoming adjacent in the 2D image, traditional 2D CNN feature extraction operators may become less effective. In response to this issue, some methods [174, 176, 395] have specifically redesigned feature extraction operators for range images. Based on the different data representation views, the view-based methods can be divided into two categories: 1) Range View, 2) BEV View.
@@ IV-A1 Range View
where denotes a 3D point and denotes the 2D map position of its projection. and denote the azimuth and elevation angle when observing the point. and are the average horizontal and vertical angle resolution between consecutive beam emitters, respectively.
VeloFCN [393] is an influential work which first introduced the projection method in cylindrical coordinates. Then, it is followed by [174, 176, 175, 173]. LaserNet [396] utilized DLA-Net [398] to obtain multi-scale features and detect 3D objects from this representation. Inspired by LaserNet, some works borrow the models in 2D object detection to handle range images, e.g. U-Net [399] is applied in [397, 400, 173], RPN [401] is employed in [173, 174] and FPN [402] is leveraged in [176]. Considering the limitations of traditional 2D CNNs in extracting features from range images, some works resort to novel operators, including range dilated convolutions [174], graph operators [395], and meta-kernel convolutions [176]. Furthermore, some works have focused on addressing occlusion and scale variation issues in range view. Specifically, these methods [175, 173] construct feature transformation structures from range view to point view and from point view to BEV view to convert range features into the BEV perspective.
@@ IV-A2 BEV View
Comparison to range view detection, BEV-based detection is more robust to occlusion and scale variation challenges. Hence, feature extraction from the range view and object detection from the bird’s eye view becomes the most practical solution to range-based 3D object detection.
The bird’s eye view representation is encoded by height, intensity and density. The point cloud is discretized into a regular 2D grid. To encode more detailed height information, the point cloud is evenly divided into slices, resulting in height maps where each grid cell stores the maximum height value of the point cloud. The intensity feature represents the reflectance value of the point within each grid cell. And the point cloud density indicates the number of points in each cell. PIXOR [403], which outputs oriented 3D object estimates decoded from pixel-wise neural network predictions, is a pioneering work in this field and followed by [404, 405, 175, 176]. These methods usually entailing three stages. First, point cloud is projected into a novel cell encoding for bird’s eye view projection. Later, both object location on the plane and its heading are estimated through a convolutional neural network originally designed for image processing. Considering scale variation and occlusion, RangeRCNN [173] and RangeIOUDet [175] introduced point view serves as a bridge from RV to BEV which provides pointwise features for models.
@@ IV-B Voxel-based 3D object detection
Voxel-based methods propose to divide the sparse point cloud and assign the distributed point cloud into regular voxels, forming the dense data representation, termed voxelization. Generally, the overall process of voxel-based methods is illustrated in Fig. 8. Compared to view-based methods, voxel-based methods leverage spatial convolution to effectively perceive 3D spatial information and achieve higher detection accuracy. Voxel-based methods face the following challenges:
High computational complexity: Compared with camera-based methods, voxel-based methods require significant memory and computation resources due to the large number of voxels used to represent the 3D space.
Spatial information loss: Due to the discrete nature of voxels, details and shape information can be lost or blurred during the voxelization process. Additionally, the limited resolution of voxels makes it challenging to accurately detect small objects.
Inconsistency in scale and density: Voxel-based methods are typically performed on voxel grids with specific scales and densities. However, due to the significant variations in object scales and point cloud densities across different scenes, making methods to adapt different scenes becomes challenging.
To overcome the aforementioned challenges, it is necessary to address the limitations of data representation, improve the network’s feature capacity and target localization accuracy, and enhance the algorithm’s understanding of complex scenes. Indeed, it is crucial for ensuring safety perception in autonomous driving. Although the optimization strategies of these methods may vary, they share common perspectives of model optimization: 1) data representation. 2) model structure.
@@ IV-B1 Data representation
Voxel-based methods first rasterize point clouds into discrete grid representations. Grid representations are closely related to accuracy, computational complexity, and memory requirements. Using too large voxel size results in significant information impairment while using too small voxel size increases the burdens of computation and memory. As shown in Fig. 9, according to the height of the z-axis, the type of grid representations can be categorized into voxel and pillar.
@@
Voxel grid process divides the 3D space into regular voxel grids with size () in the x, y, and z directions, respectively. Only non-empty voxel units that contain points are stored and used for feature extraction. However, due to the sparse distribution of point clouds, the majority of voxel units are empty. As a pioneering work in voxel-based methods [112, 115, 113, 109, 129, 130, 105, 151, 143, 101, 102, 103, 125, 251, 182, 104, 121, 122, 136, 230, 114, 108, 131],VoxelNet [108] proposes a novel voxel feature encoding (VFE) layer to extract features from the points inside a voxel cell. Theb, following works [113, 149, 114, 104, 103, 125, 161] have extended the VoxelNet network by adopting similar voxel encoding approaches. Existing methods often perform local partitioning and feature extraction uniformly across all positions in the point cloud. This approach leads to limitations in the receptive field for distant regions and information truncation. Therefore, some works have proposed different approaches to voxel partitioning: 1) Different coordinate systems. Some approaches have reexamined voxel partitioning from different coordinate system perspectives, e.g. [200, 153] from cylindrical and [132] from spherical coordinate systems. Sphereformer [132] facilitates the aggregation of information from sparsely distant points by dividing the 3D space into multiple non-overlapping radial windows using spherical coordinates (, , ), thereby enhancing information integration from dense point regions. 2) Multi-scale voxels. Some works generate voxels of different scales [115, 150] or use reconfigurable voxels [155], e.g. HVNet [150] proposes a hybrid voxel network which integrats different scales in the point-level voxel feature encoder (VFE). In addition, there are two series of approaches trying to incorporate additional information for voxel fusion. 1) Additional temporal information. Some methods [122, 126, 406, 407, 408, 409, 410] integrate point cloud data from multiple time steps to obtain global environmental information, effectively mitigating the effects of objects blocking each other, and providing a denser representation that captures more detailed spatial information. 2) Additional spatial prior information. Some works encode density information [120, 411, 412, 413] or voxel centroid information [120] in point cloud processing.
Pillar
@@
Pillars can be considered as a special form of voxels. Specifically, point clouds are discretized into a grid uniformly distributed on the x-y plane, creating a set of pillars without binning along the z-axis. Pillar features can be aggregated from points through a PointNet [414] and then scattered back to construct a 2D BEV image for feature extraction. As the pioneering work in this series [183, 118, 135, 119, 181, 123, 156], PointPillar [118] firstly introduces pillar representation. Followed works have extended the ideas from 2D detection to PointPillars. PillarNet [183] adopts the ‘encoder-neck-head’ detection architecture to enhance the performance of pillar-based methods. SWFormer [135] and ESS [123] draw inspiration from the swin transformer [352] and apply a hierarchical window mechanism to pseudo-images, enabling the network to maintain a global receptive field while achieving faster inference speed. PillarNext [119] integrates a series of mature 2D detection techniques and achieves performance comparable to voxel-based methods.

### v2

@@ III-D1 Accuracy
Accuracy is a focal point of interest in most research articles and reviews and is indeed of paramount importance. While accuracy can be reflected through AP (average precision), considering AP alone for comparison may not provide a comprehensive view, as different methodologies may exhibit substantial differences due to differing paradigms.
As shown in Fig. 3 (a), we selected ten representative methods (including classic and latest research) for comparison, and it is evident that there are significant metric disparities between monocular 3D object detection [179, 183, 13, 23, 29, 10, 28, 184] and stereo-based 3D object detection [286, 216, 231, 14, 236, 378, 237, 17, 234]. The current scenario indicates that the accuracy of monocular 3D object detection is far lower than that of stereo-based 3D object detection. Stereo-based 3D object detection leverages the capture of images from two different perspectives of the same scene to obtain depth information. The greater the baseline between cameras, the wider the range of depth information captured. As shown in Fig. 3 (b), there were monocular 3D object detection methods [379, 9, 25, 27, 24] on the nuScenes dataset [134], but no related research on stereo-based 3D object detection.
Starting from 2021, monocular methods have gradually been supplanted by multi-view (bird’s-eye-view perception) 3D object detection methods [30, 31, 32, 16, 282, 273, 15, 272, 279, 280], leading to a significant improvement in mAP. The emergence of the novel bird’s-eye-view paradigm and the increase in sensor quantity have substantially impacted mAP. It can be observed that initially, the disparity between DD3D [27] and DETR3D [30] is not prominent, but with the continuous enhancement of multi-view 3D object detection, particularly with the advent of novel works such as Far3D [280], the gap has widened. In other words, camera-only 3D object detection methods on multi-camera datasets like nuScenes [134] are predominantly based on bird’s-eye-view perception. If we consider accuracy solely from this single dimension, the increase in sensor quantity has significantly improved accuracy metrics (including mAP, NDS, AP, etc.).
@@ III-D2 Latency
In 3D object detection, latency (frames er second, FPS) and accuracy are critical metrics for evaluating algorithm performance [386]. As shown in Table IV, Monocular-based 3D object detection, which relies on data from a single camera, typically achieves higher FPS due to lower computational requirements. However, its accuracy is often inferior to stereo or multi-view systems due to the absence of depth information. Stereo-based detection, leveraging disparity information from dual cameras, enhances depth estimation accuracy but introduces greater computational complexity, potentially reducing FPS. Multi-view detection provides richer scene information and improved accuracy but demands extensive data processing, computational power, and algorithmic optimization for reasonable FPS levels. Notably, the nuScenes dataset lacks representation of stereo-based methods, with the monocular method FCOS3D[24] standing out as emblematic, introduced in 2021. Over time, multi-view 3D object detection has rapidly evolved in terms of accuracy and latency. In practice, real-time performance is also an important consideration when deploying a robust 3D object detection system. For example, ER3D[387] takes stereo images as input and predicts 3D bounding boxes, which leverages a fast but inaccurate method of semi-global matching for depth estimation. Li et al. [388] propose a lightweight Pseudo-LiDAR 3D detection system that achieves high accuracy and responsiveness. RTS3D [231] proposes a novel framework for faster and more accurate 3D object detection using stereo images. FastFusion [389], a three-stage stereo-LiDAR deep fusion scheme, integrates LiDAR priors into each step of the classical stereo-matching taxonomy, thereby gaining high-precision dense depth sensing in real-time. In conclusion, achieving safe autonomous driving necessitates balancing latency and accuracy in 3D object detection algorithms. While monocular detection is faster, it lacks precision. Stereo and multi-view methods are accurate but slower. Future research should focus on maintaining high precision while emphasizing increased FPS and reduced latency to meet the dual requirements of real-time responsiveness and safety in autonomous driving.
@@ III-D3 Robustness
Robustness constitutes a pivotal factor in the safety perception of autonomous driving, representing a topic of significant attention that has been previously overlooked in comprehensive reviews. In the current meticulously designed clean datasets and benchmarks, such as KITTI [133], nuScenes [134], and Waymo [135], this aspect is not commonly addressed. Presently, research works[124, 125, 390, 147, 148, 391, 392] like RoboBEV[124], Robo3D[148] on 3D object detection incorporate considerations of robustness, exemplified by factors such as sensor misses, as illustrated in Fig. 4. They have adopted a methodology involving the introduction of disturbances into datasets relevant to 3D object detection to assess robustness. This includes introducing various types of noise, such as variations in weather conditions, sensor malfunctions, motion disturbances, and object-related perturbations, aimed at unraveling the distinct impacts of different noise sources on the model. Typically, most papers investigating robustness conduct evaluations by introducing noise to the validation sets of clean datasets, such as KITTI [133], nuScenes [134], and Waymo [135]. Additionally, we highlight findings from Ref. [125], where KITTI-C[125] and nuScenes-C[125] are emphasized as examples to illustrate the results of Camera-Only 3D object detection methods. Tables VII and VIII provide an overall comparison, revealing that, in general, Camera-Only methods are less robust compared to LiDAR-Only and multi-modal fusion methods. They are highly susceptible to various types of noise. In KITTI-C, three representative works—SMOKE[179], PGD[25], and ImVoxelNet[393]—show consistently lower overall performance and reduced robustness to noise. In nuScenes-C, noteworthy methods such as DETR3D[30] and BEVFormer[16] exhibit greater robustness compared to FCOS3D [24] and PGD[25], suggesting that as the number of sensors increases, overall robustness improves. In conclusion, future Camera-Only methods need to consider not only cost and accuracy metrics (mAP, NDS, etc.) but also factors related to safety perception and robustness. Our analysis aims to provide valuable insights for the safety of future autonomous driving systems.
| Corruptions | LiDAR-Only | Camera-Only | Multi-modal | |||||||||||
| SECOND † | PointPillars † | PointRCNN † | PV-RCNN † | Part-A2 † | 3DSSD † | SMOKE † | PGD † | ImVoxelNet † | EPNet † | Focals Conv † | LoGoNet * | VirConv-S * | ||
@@
| Corruptions | LiDAR-Only | Camera-Only | Multi-modal | |||||||||
| PointPillars† | SSN† | CenterPoint† | FCOS3D† | PGD† | DETR3D† | BEVFormer† | FUTR3D† | TransFusion† | BEVFusion† | D.I.* | ||
@@
IV LiDAR-only 3D Object Detection
LiDAR-only methods capture precise 3D information, leading to higher detection accuracy and robustness, particularly in extreme weather conditions [125]. Because in comparison to optical radiation, the laser beams emitted by LiDAR systems can penetrate certain weather disturbances, such as raindrops and haze, with slight interference. However, the high cost of LiDAR remains one of the main barriers to large-scale adoption of LiDAR-only methods. Generally, as shown in Fig. 5, LiDAR-only methods can be categorized into four types: (1) view-based 3D object detection, (2) voxel-based 3D object detection, (3) point-based 3D object detection, (4) point-voxel-based 3D object detection. In contrast to previous reviews [4, 126, 123, 130, 131], our survey extends beyond the conventional classifications of LiDAR-only methods. We adopt a more foundational idea to class LiDAR-only methods based on their core data representations (BEV, Voxel, Pillars.) and underlying model structure (CNNs, Transformers, PointNet). We provide a comprehensive understanding of the technological paradigms at LiDAR-only methods, analyzing and classifying these systems from a more essential, technical lineage perspective.
@@ IV-A View-based 3D object detection
View-based methods transform point clouds into pseudo-images using BEV and range views. Based on the different data representation views, the view-based methods can be divided into two categories: 1) Range View, 2) BEV View. In these representations, each pixel contains 3D spatial information rather than RGB values. Due to the dense representation of pseudo-images, traditional or specialized 2D convolutions can be seamlessly applied to range images, making the feature extraction process highly efficient. However, compared to other LiDAR-only methods, detection using range views is more susceptible to occlusion and scale variations.
@@ IV-A1 Range View
where denotes a 3D point and denotes the 2D map position of its projection. and denote the azimuth and elevation angle when observing the point. and are the average horizontal and vertical angle resolution between consecutive beam emitters, respectively. VeloFCN [396] is an influential work that first introduces the projection method in cylindrical coordinates. It has then followed by [88, 90, 89, 87]. LaserNet [394] utilizes DLA-Net [397] to obtain multi-scale features and detect 3D objects from this representation. Inspired by LaserNet, some works have borrowed models from 2D object detection to handle range images. For example, U-Net [398] is applied in [395, 399, 87], RPN [400] is employed in [87, 88], and FPN [401] is leveraged in [90]. Considering the limitations of traditional 2D CNNs in extracting features from range images, some works have resorted to novel operators, including range dilated convolutions [88], graph operators [402], and meta-kernel convolutions [90]. Furthermore, some works have focused on addressing issues of occlusion and scale variation in range view. Specifically, these methods [89, 87] construct feature transformation structures from the range view to the point view and from the point view to the BEV (Bird’s Eye View) perspective to convert range features into BEV perspective.
@@ IV-A2 BEV View
Comparison to range view detection, BEV-based detection is more robust to occlusion and scale variation challenges. Hence, feature extraction from the range view and object detection from the BEV become the most practical solution for range-based 3D object detection. The BEV representation is encoded by height, intensity, and density. Point clouds are discretized into a regular 2D grid. To encode more detailed height information, point clouds are evenly divided into slices, resulting in height maps where each grid cell stores the maximum height value of the point clouds. The intensity feature represents the reflectance value of the point within each grid cell, and the point cloud density indicates the number of points in each cell. PIXOR [291], which outputs oriented 3D object estimates decoded from pixel-wise neural network predictions, is a pioneering work in this field, followed by [293, 403, 89, 90]. These methods usually entail three stages. First, point clouds are projected into a novel cell encoding for BEV projection. Next, both the object’s location on the plane and its heading are estimated through a convolutional neural network originally designed for image processing. Considering scale variation and occlusion, RangeRCNN [87] and RangeIOUDet [89] introduce a point view that serves as a bridge from RV to BEV, which provides pointwise features for the models.
@@ IV-B Voxel-based 3D object detection
Voxel-based methods segment sparse point clouds into regular voxels, achieving a dense representation through voxelization. Despite spatial convolution enhancing 3D information perception, challenges persist in achieving high detection accuracy. These challenges include 1) high computational complexity, which demands substantial memory and computational resources due to the numerous voxels representing 3D space, 2) spatial information loss that occurs during voxelization, leading to difficulties in accurately detecting small objects, 3) inconsistencies in scale and density, inherent to specific voxel grids, which pose challenges in adapting to diverse scenes with varying object scales and point cloud densities. Overcoming these challenges requires addressing limitations in data representation, enhancing network feature capacity, improving object localization accuracy, and enhancing the model’s understanding of complex scenes. Ensuring safety perception in autonomous driving is crucial, and despite varying optimization strategies, these methods converge on common perspectives of model optimization, focusing on 1) data representation and 2) model structure.
@@ IV-B1 Data representation
Voxel-based methods first rasterize point clouds into discrete grid representations. Grid representations are closely related to accuracy, computational complexity, and memory requirements. Using a voxel size that is too large results in significant information loss, while using a voxel size that is too small increases the burdens of computation and memory. As shown in Fig. 5 (b), according to the height along the z-axis, the types of grid representations can be categorized into voxels and pillars.
@@
Voxel process divides the 3D space into regular voxel grids with size () in the x, y, and z directions, respectively. Only non-empty voxel units that contain points are stored and used for feature extraction. However, due to the sparse distribution of point clouds, the majority of voxel units are empty. As a pioneering work in voxel-based methods [46, 48, 47, 43, 60, 61, 40, 77, 69, 36, 37, 38, 57, 92, 39, 54, 65, 108, 42], VoxelNet [42] proposes a novel voxel feature encoding (VFE) layer to extract features from the points inside a voxel cell. Then, following works [47, 75, 39, 38, 57, 80] have extended the VoxelNet network by adopting similar voxel encoding approaches. Existing methods often perform local partitioning and feature extraction uniformly across all positions in the point cloud. This approach limits the receptive field for distant regions and information truncation. Therefore, some works have proposed different approaches to voxel partitioning: 1) Different coordinate systems: some approaches have reexamined voxel partitioning from different coordinate system perspectives, e.g. [404, 78] from cylindrical and [62] from spherical coordinate systems. Sphereformer [62] facilitates the aggregation of information from sparsely distant points by dividing the 3D space into multiple non-overlapping radial windows using spherical coordinates (, , ), thereby enhancing information integration from dense point regions. 2) Multi-scale voxels: some works generate voxels of different scales [48, 76] or use reconfigurable voxels [79], e.g., HVNet [76] proposes a hybrid voxel network which integrates different scales in the point-level voxel feature encoder (VFE).
Pillars
@@
Pillars can be considered a special form of voxels. Specifically, point clouds are discretized into a grid uniformly distributed on the x-y plane without binning along the z-axis. Pillar features can be aggregated from points through a PointNet [405] and then scattered back to construct a 2D BEV image for feature extraction. As the pioneering work in this series [93, 51, 64, 52, 91, 55], PointPillar [51] first introduces the pillar representation. Following works have extended the ideas from 2D detection to PointPillars. PillarNet [93] adopts the ’encoder-neck-head’ detection architecture to enhance the performance of pillar-based methods. SWFormer [64] and ESS [55] draw inspiration from the Swin Transformer [371] and apply a hierarchical window mechanism to pseudo-images, thereby enabling the network to maintain a global receptive field. PillarNeXt [52] integrates a series of mature 2D detection techniques and achieves performance comparable to voxel-based methods.